what is the gauge pressure at the bottom of a freshwater pool that is 12.5 ft deep?


1
Expert's answer
2021-07-26T17:11:01-0400

"P=\rho gh"

"P=10^3Kg/m^3×9.8×(\frac {12.5×12×2.54}{100}m)"

"P=37,338Pa"

"P=37,338\u00d70.000145"

"P=5.4(psi)"